How much energy is needed to change 5g of ice at 0 degrees Celsius to water at 20 degrees Celsius, given that specific heat capacity of water is 4. 2 J / g degrees Celsius.


1
Expert's answer
2020-07-20T15:33:17-0400

We can find the amount of energy from the formula:


"Q = mL_f + mc \Delta T,""Q = 5 \ g \cdot 336 \ \dfrac{J}{g} + 5 \ g \cdot 4.2 \ \dfrac{J}{g \cdot \! ^{\circ}C} \cdot 20 \ ^{\circ}C = 2100 \ J."

Answer:

"Q = 2100 \ J."